OpenStudy (anonymous):

rationalize the denominator: 2/(sqrt3 - sqrt2)

OpenStudy (anonymous):

\[2/\sqrt{3}-\sqrt{2}\]

OpenStudy (ajprincess):

multiply by \(\sqrt3+\sqrt2\) up and down. tel me vat u get.

OpenStudy (anonymous):

\[2\sqrt{3}+2\sqrt{2}\]

OpenStudy (anonymous):

\[(\sqrt{a} - \sqrt{b})(\sqrt{a} + \sqrt{b}) = a - b\]

OpenStudy (ajprincess):

vat abt the denominator

OpenStudy (anonymous):

doesnt it become 1?

OpenStudy (anonymous):

Yes you are right...

OpenStudy (anonymous):

so you're just left with the numerator?

OpenStudy (ajprincess):

yes.